# Small maximal clusters are very unlikely in critical random graphs

### v1

@@ 1.4 Notation and structure of the paper
$\mathbb{P}\big{(}\exists t\leq(t_{i}-t_{i-1})\wedge b:h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\mid\mathcal{F}_{i-1}\big{)}$ $+\mathbb{P}\big{(}\exists b<t\leq(t_{i}-t_{i-1})\wedge T:\sum_{i=1}^{t}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq t^{\gamma}\mid\mathcal{F}_{i-1}\big{)}$
$\leq\frac{2b^{2}}{n}+\mathbb{P}\big{(}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\mid\mathcal{F}_{i-1}\big{)}$
$\leq\frac{2b^{2}}{n}+\exp\big{(}-b^{\gamma}+cT\ell n^{-1}\big{)}+\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{i-1}).$ (65)
There remains to bound the (conditional) probability of $\mathcal{B}^{c}$, given $\mathcal{F}_{i-1}$. This is easily done (here we follows closely Lemma $4.6$ in [13]). Note that
$\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{i-1})\leq\mathbb{P}\big{(}\exists t\leq(t_{i}-t_{i-1})\wedge T:d_{i}+\sum_{k=1}^{t}D^{i}_{k}>\ell\big{)},$ (66)
where $D^{i}_{k}\coloneqq\mathbbm{1}_{R^{i}_{k}}(d-1)-1$ (the upper bound is justified by the fact that, for $t\leq t_{i}-t_{i-1}$, we can write $|\mathcal{A}^{i}_{t}|=d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}$ and each $\eta^{i}_{k}$ is at most $D^{i}_{k}$; moreover, the $R^{i}_{k}$ do not depend on $\mathcal{F}_{i-1}$, as they only pertain to the percolation process on $\mathbb{G}^{\prime}(n,d)$). Recalling that $R^{i}_{k}$ occurs with probability $p=(1+\lambda n^{-1/3})/(d-1)$ (being the event that the edge revealed at step $k$ is retained), we see that $\sum_{k=1}^{t}D^{i}_{k}$ is a sub-martingale for $\lambda>0$ (an assumption which can be made without loss of generality); an application of lemma 1.2 (i.e. Doob’s inequality) as in Lemma $4.6$ of [13] yields that the probability on the right-hand side of (66) is at most
@@
for positive constants $C,C^{\prime}>0$. Setting for instance $b\coloneqq\lceil A^{1/5}n^{1/3}\rceil$ and $\ell\coloneqq n^{1/3+\gamma/3}A$ we see that $T\ell n^{-1}\ll b^{\gamma}$ as well as $\ell^{2}/T\gg\ell/n^{1/3}$. Thus, the exponential term in (2.3.1) is at most $e^{-cb^{\gamma}}$, whereas the expression on the right-hand side of (67) is at most $e^{-c\ell^{2}/T}$. Both these terms are, for all large enough $n$, at most $b^{2}/n\asymp A^{2/5}/n^{1/3}$, which in turn is $\ll A^{1/2}/n^{1/3}\asymp T^{-1/2}$, concluding the proof of the lemma. ∎
@@ 3 Boosting the martingale method of Nachmias and Peres
The proof of Theorem 1.1 rely on random walk estimates which are easily applicable provided that the (conditional) distribution of the number of unseen vertices/stubs becoming active at any step of the underlying exploration process is ‘simple enough’. When the above-mentioned conditional law is more complicated, an analysis like the one we provided in Section 2 becomes more involved. To this end, the goal of this section is to show how to optimize the (very robust) martingale argument of Nachmias and Peres [28, 27] in order to obtain stretched-exponential upper bounds for the probability of observing an unusually small maximal cluster in critical random graphs. With our adjustments, we can show that the above probability is at most $\exp(-A^{3/5})$ in other two (critical) models (which are introduced below), thus considerably improving upon the existing polynomial bounds available in the literature for such random graphs; this is the content of Theorem 3.1 below. (We remark that the same methodology used to establish the bounds of Theorem 3.1 applies to the models treated in Theorem 1.1; however, since we already derived better bounds for those random graphs in Section 2, we won’t treat them again here.)
For the random graphs treated in Theorem 3.1 we are content with the stretched bounds given here, mostly because we believe that the optimised version of the martingale argument (described below) is interesting on its own and provides a good alternative to the more precise methodology of Section 2. However, we believe that by adapting the same argument used to analyse the near-critical $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$ models, it should be possible to show that the probability of observing an unusually small maximal components in these other (critical) models is also at most of order $\exp(-A^{3/2})$. However, such an adaptation requires some work, since the (conditional) distributions of the random variables $\eta_{i}$ are now more complicated: they are still binomials, but now also the success probabilities are random (and consequently the random walk estimates used to treat the models of Theorem 1.1 requires some work to be implemented).
Before stating Theorem 3.1, we first need to introduce the critical random graphs considered there.
@@
As we pointed out at the start of this section, the same method used to establish Theorem 3.1 can also be used to prove similar bounds for the critical random graphs $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$; however, we have already established stronger estimates in Section 2 and hence there is no reason to use the optimised martingale argument to obtain bounds of the type $\exp(-cA^{3/5})$ for these two models.
We continue by providing a detailed, but informal explanation on how to optimize the martingale argument introduced in [28, 27] to obtain the stretched exponential upper bounds displayed in Theorem 3.1; we do so in terms of the critical $\mathbb{G}(n,p)$ model, but the same strategy works for the critical random graphs $\mathbb{G}(n,k,p),\mathbb{G}(n,\beta,\lambda)$ (and $\mathbb{G}(n,d,p)$).
@@ 3.1 Heuristic derivation of the stretched exponential bounds
Here we show how the original argument of [28] can be boosted to show that $|\mathcal{C}_{\max}(\mathbb{G}(n,1/n))|\geq n^{2/3}/A$ with probability at least of order $1-\exp(-A^{3/5})$; we provide the explanation in terms of the $\mathbb{G}(n,1/n)$ model for simplicity of exposition, but the same argument applies to the other models considered in this work.
@@
To obtain an exponential (upper) bound for the first probability in (73) instead, we proceed as follows. Let $\tau$ be the minimum between the first time $t\in\mathbb{N}$ at which the number of active vertices $Y_{t}$ goes above $h\delta$ and $T=mh^{2}$. Denote by $\eta_{t}$ the number of unseen vertices which are added to the set of active nodes at step $t$. We start by noticing that the process $Y_{\tau+s}$ (started at time $\tau$) must reach zero in less than $n^{2/3}/A$ steps, otherwise we would discover a positive excursion consisting of more than $n^{2/3}/A$ steps. Hence there is a first time $t<n^{2/3}/A$ at which $Y_{\tau+t}=0$. Observe that at time $\tau+i$ we add $\eta_{\tau+i}$ vertices to the set of active nodes and remove one vertex from such a set (as long as the set of active nodes remains non-empty); thus, since the number of active vertices stays positive from time $\tau$ until time $\tau+t$, we can write $Y_{\tau+s}=Y_{\tau}+\sum_{i=1}^{s}(\eta_{\tau+i}-1)$ for $s\leq t$. But we know that $Y_{\tau}\geq h\delta$ (and $Y_{\tau+t}=0$), whence we arrive at $\sum_{i=1}^{t}(\eta_{\tau+i}-1)\leq-h\delta$. Conditional on the history of the exploration process until step $\tau+i-1$, the random variable $\eta_{\tau+i}$ has the $\text{Bin}(n-\tau-i+1-Y_{\tau+i-1},p)$ distribution (with $n-\tau-i+1-Y_{\tau+i-1}$ being the number of unseen vertices available at time $\tau+i$). Since the number of active vertices is never ‘too large’, to continue our derivation we approximate $\text{Bin}(n-\tau-i-1-Y_{\tau+i-1},p)\approx\text{Bin}(n-\tau-i,p)$. Then, roughly, $\sum_{i=1}^{t}(\eta_{\tau+i}-1)\leq-h\delta$ occurs if, and only if, $\sum_{i=1}^{t}(\text{Bin}_{i}(n-\tau-i,p)-1)\leq-h\delta$ happens. Since $\tau\leq mh^{2}$ and $t\leq n^{2/3}/A$, the latter event ‘implies’ that $\sum_{i=1}^{t}(\text{Bin}_{i}(n-mh^{2}-n^{2/3}/A,p)-1)\leq-h\delta$ occurs. Moreover, using the fact that binomial random variables concentrate around their mean (whence in particular $\sum_{i=1}^{t}\text{Bin}_{i}(T+n^{2/3}/A,p)\approx(tmh^{2})/n+t/(n^{1/3}A)$, where we used that $p=1/n$ and $t<n^{2/3}/A$), we conclude that the probability of $\sum_{i=1}^{t}(\text{Bin}_{i}(n-mh^{2}-n^{2/3}/A,p)-1)\leq-h\delta$ happening at some $t<n^{2/3}/A$ is (very) roughly the same as the probability of $\sum_{i=1}^{t}(\text{Bin}_{i}(n,p)-1)\leq-h\delta+(tmh^{2})/n+t/(n^{1/3}A)$ occurring at some $t<n^{2/3}/A$. Now notice that, since $t<n^{2/3}/A$, we can bound
@@
and hence the last probability is at most the probability of seeing a time $t<n^{2/3}/A$ at which $\sum_{i=1}^{t}(\text{Bin}_{i}(n,p)-1)\leq-h\delta+(mh^{2})/(n^{1/3}A)+n^{1/3}/A^{2}$, which is small only if both $(mh^{2})/(n^{1/3}A)$ and $n^{1/3}/A^{2}$ are sufficiently smaller than $h\delta$, say smaller than $(h\delta)/3$. In this case then, we would we bounding from above the probability that a mean-zero random walk (with supposedly iid, well-behaved) $\text{Bin}_{i}(n,p)-1$ increments goes below $\asymp-h$ in less than $n^{2/3}/A$ steps. An application of Doob’s submartingale inequality yields that this occurs with probability at most $\lesssim\exp(-(h^{2}A)/n^{2/3})$. Combining this estimate with our earlier bound of order $e^{-m}$ we see that the sum of the two terms in (73) is at most of order
@@ 3.3 The $\mathbb{G}(n,\beta,\lambda)$ model
Then, arguing as we did for the random graphs $\mathbb{G}_{n,p}$ and $\mathbb{G}_{n,d,p}$, we take $1\ll h=h(n)\in\mathbb{N}$, $m\in\mathbb{N}$ (possibly dependent on $n$), and denote by $\delta\in(0,1)$ a constant that we specify later. Then, setting $T\coloneqq\lceil n^{2/3}/A\rceil$, we bound from above the probability on the right-hand side of (106) by

### v3

@@ 1.4 Some useful tools
$\mathbb{P}\big{(}\exists t\leq(t_{i}-t_{i-1})\wedge b:h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\mid\mathcal{F}_{t_{i-1}}\big{)}$ $+\mathbb{P}\big{(}\exists b<t\leq(t_{i}-t_{i-1})\wedge T:\sum_{i=1}^{t}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq t^{\gamma}\mid\mathcal{F}_{t_{i-1}}\big{)}$
$\leq\frac{2b^{2}}{n}+\mathbb{P}\big{(}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\mid\mathcal{F}_{t_{i-1}}\big{)}$
$\leq\frac{2b^{2}}{n}+\exp\big{(}-b^{\gamma}+cT\ell n^{-1}\big{)}+\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{t_{i-1}}).$ (65)
There remains to bound the (conditional) probability of $\mathcal{B}^{c}$, given $\mathcal{F}_{t_{i-1}}$. This is easily done (here we follows closely Lemma $4.6$ in [13]). Note that
$\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{t_{i-1}})\leq\mathbb{P}\big{(}\exists t\leq(t_{i}-t_{i-1})\wedge T:d_{i}+\sum_{k=1}^{t}D^{i}_{k}>\ell\big{)},$ (66)
where $D^{i}_{k}\coloneqq\mathbbm{1}_{R^{i}_{k}}(d-1)-1$ (the upper bound is justified by the fact that, for $t\leq t_{i}-t_{i-1}$, we can write $|\mathcal{A}^{i}_{t}|=d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}$ and each $\eta^{i}_{k}$ is at most $D^{i}_{k}$; moreover, the events $R^{i}_{k}$ do not depend on $\mathcal{F}_{t_{i-1}}$, as they only pertain to the percolation process on $\mathbb{G}^{\prime}(n,d)$). Recalling that $R^{i}_{k}$ occurs with probability $p=(1+\lambda n^{-1/3})/(d-1)$ (being the event that the edge revealed at step $t_{i-1}+k$ is retained), we see that $\sum_{k=1}^{t}D^{i}_{k}$ is a sub-martingale for $\lambda>0$ (an assumption which can be made without loss of generality); an application of lemma 1.2 (i.e. Doob’s inequality) as in Lemma $4.6$ of [13] yields that the probability on the right-hand side of (66) is at most
@@
for positive constants $C,C^{\prime}>0$. Setting for instance $b\coloneqq\lceil A^{1/5}n^{1/3}\rceil$ and $\ell\coloneqq n^{1/3+\gamma/3}A$ we see that $T\ell n^{-1}\ll b^{\gamma}$ as well as $\ell^{2}/T\gg\ell/n^{1/3}$. Thus, the exponential term in (2.3.1) is at most $e^{-cb^{\gamma}}$, whereas the expression on the right-hand side of (67) is at most $e^{-c\ell^{2}/T}$. Both these terms are, for all large enough $n$, at most $b^{2}/n=\Theta(A^{2/5}/n^{1/3})$, which in turn is $\ll A^{1/2}/n^{1/3}=\Theta(T^{-1/2})$, concluding the proof of the lemma. ∎
@@ 3 Boosting the martingale method of Nachmias and Peres
The proof of Theorem 1.1 rely on random walk estimates which are easily applicable provided that the (conditional) distribution of the number of unseen vertices/stubs becoming active at any step of the underlying exploration process is ‘simple enough’. When the above-mentioned conditional law is more complicated, an analysis like the one we provided in Section 2 becomes a bit more involved. To this end, the goal of this section is to show how to optimize the (very robust) martingale argument of Nachmias and Peres [28, 27] in order to obtain stretched-exponential upper bounds for the probability of observing an unusually small maximal cluster in critical random graphs. With our adjustments, we can show that the above probability is at most $\exp(-A^{3/5})$ in other two (critical) models (which are introduced below), thus considerably improving upon the existing polynomial bounds available in the literature for such random graphs; this is the content of Theorem 3.1 below. We remark that the same methodology used to establish the bounds of Theorem 3.1 applies to the models treated in Theorem 1.1; however, since we already derived better bounds for those random graphs in Section 2, we won’t treat them again here.
For the random graphs treated in Theorem 3.1 we are content with the stretched bounds given here, mostly because we believe that the optimized version of the martingale argument (described below) is interesting on its own and provides a good alternative to the more precise methodology of Section 2. However, we believe that by adapting the same argument used to analyze the near-critical $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$ models, it should be possible to show that the probability of observing an unusually small maximal components in these other (critical) models is also at most $\exp(-\Omega(A^{3/2}))$. However, such an adaptation requires some work, since the (conditional) distributions of the random variables $\eta_{i}$ are now more complicated: they are still binomials, but now also the success probabilities are random (and consequently the random walk estimates used to treat the models of Theorem 1.1 requires some work to be implemented). Before stating Theorem 3.1, we first need to introduce the critical random graphs considered there.
@@
As we pointed out at the start of this section, the same method used to establish Theorem 3.1 can also be used to prove similar bounds for the critical random graphs $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$; however, we have already established stronger estimates in Section 2 and hence there is no reason to use the optimized martingale argument to obtain bounds of the type $\exp(-cA^{3/5})$ for these two models.
We continue by providing a detailed, but still informal explanation on how to optimize the martingale argument introduced in [28, 27] to obtain the stretched exponential upper bounds displayed in Theorem 3.1; we do so in terms of the critical $\mathbb{G}(n,p)$ model, but the same strategy works for the critical random graphs $\mathbb{G}(n,k,p),\mathbb{G}(n,\beta,\lambda)$ and $\mathbb{G}(n,d,p)$.
@@ 3.1 Heuristic derivation of the stretched exponential bounds
Here we show how the original argument of [28] can be boosted to show that $|\mathcal{C}_{\max}(\mathbb{G}(n,1/n))|\geq n^{2/3}/A$ with probability at least $1-\exp(-\Omega(A^{3/5}))$; we provide the explanation in terms of the $\mathbb{G}(n,1/n)$ model for simplicity of exposition, but the same argument applies to the other models considered in this work.
@@
To obtain an exponential (upper) bound for the first probability in (73) instead, we proceed as follows. Let $\tau$ be the minimum between the first time $t\in\mathbb{N}$ at which the number of active vertices $Y_{t}$ goes above $h\delta$ and $T=mh^{2}$. Denote by $\eta_{t}$ the number of unseen vertices which are added to the set of active nodes at step $t$. We start by noticing that the process $Y_{\tau+s}$ (started at time $\tau$) must reach zero in less than $n^{2/3}/A$ steps, otherwise we would discover a positive excursion consisting of more than $n^{2/3}/A$ steps. Hence there is a first time $t<n^{2/3}/A$ at which $Y_{\tau+t}=0$. Observe that at time $\tau+i$ we add $\eta_{\tau+i}$ vertices to the set of active nodes and remove one vertex from such a set (as long as the set of active nodes remains non-empty); thus, since the number of active vertices stays positive from time $\tau$ until time $\tau+t$, we can write $Y_{\tau+s}=Y_{\tau}+\sum_{i=1}^{s}(\eta_{\tau+i}-1)$ for $s\leq t$. But we know that $Y_{\tau}\geq h\delta$ (and $Y_{\tau+t}=0$), whence we arrive at $\sum_{i=1}^{t}(\eta_{\tau+i}-1)\leq-h\delta$. Conditional on the history of the exploration process until step $\tau+i-1$, the random variable $\eta_{\tau+i}$ has the $\text{Bin}(n-\tau-i+1-Y_{\tau+i-1},p)$ distribution (with $n-\tau-i+1-Y_{\tau+i-1}$ being the number of unseen vertices available at time $\tau+i$). Since the number of active vertices is never ‘too large’, to continue our derivation we approximate
$\text{Bin}(n-\tau-i-1-Y_{\tau+i-1},p)\approx\text{Bin}(n-\tau-i,p).$
Then, roughly,
$\sum_{i=1}^{t}(\eta_{\tau+i}-1)\leq-h\delta\text{ occurs if, and only if, }\sum_{i=1}^{t}(\text{Bin}_{i}(n-\tau-i,p)-1)\leq-h\delta\text{ happens}.$
Since $\tau\leq mh^{2}$ and $t\leq n^{2/3}/A$, the latter event ‘implies’ that
$\sum_{i=1}^{t}(\text{Bin}_{i}(n-mh^{2}-n^{2/3}/A,p)-1)\leq-h\delta\text{ occurs}.$ (74)
Moreover, using the fact that binomial random variables concentrate around their mean (whence in particular
$\sum_{i=1}^{t}\text{Bin}_{i}(T+n^{2/3}/A,p)\approx(tmh^{2})/n+t/(n^{1/3}A),$
where we used that $p=1/n$ and $t<n^{2/3}/A$), we conclude that the probability of (74) happening at some $t<n^{2/3}/A$ is (very) roughly the same as the probability of
$\sum_{i=1}^{t}(\text{Bin}_{i}(n,p)-1)\leq-h\delta+(tmh^{2})/n+t/(n^{1/3}A)$ (75)
occurring at some $t<n^{2/3}/A$. Now notice that, since $t<n^{2/3}/A$, we can bound
@@
and hence the probability in (75) is at most the probability of seeing a time $t<n^{2/3}/A$ at which
$\sum_{i=1}^{t}(\text{Bin}_{i}(n,p)-1)\leq-h\delta+(mh^{2})/(n^{1/3}A)+n^{1/3}/A^{2},$
which is small only if both $(mh^{2})/(n^{1/3}A)$ and $n^{1/3}/A^{2}$ are sufficiently smaller than $h\delta$, say smaller than $(h\delta)/3$. In this case then, we would we bounding from above the probability that a mean-zero random walk (with supposedly iid, well-behaved) $\text{Bin}_{i}(n,p)-1$ increments goes below $\asymp-h$ in less than $n^{2/3}/A$ steps. An application of Doob’s submartingale inequality yields that this occurs with probability at most $\lesssim\exp(-(h^{2}A)/n^{2/3})$. Combining this estimate with our earlier bound of order $e^{-m}$ we see that the sum of the two terms in (73) is at most of order
@@ 3.3 The $\mathbb{G}(n,\beta,\lambda)$ model
Then, arguing as we did for the random graphs $\mathbb{G}(n,p)$ and $\mathbb{G}_{n,d,p}$, we take $1\ll h=h(n)\in\mathbb{N}$, $m\in\mathbb{N}$ (possibly dependent on $n$), and denote by $\delta\in(0,1)$ a constant that we specify later. Then, setting $T\coloneqq\lceil n^{2/3}/A\rceil$, we bound from above the probability on the right-hand side of (108) by